OpenStudy (anonymous):

PLEASE HELP?!?! Consider the line y=-9x+3 Find the equation of the line that is perpendicular to this line and passes through the point (6,3). Find the equation of the line that is parallel to this line and passes through the point (6,3) *I know the slope for the perpendicular is 1/9 and for the parallel its -9. But I cant figure out b...

OpenStudy (anonymous):

You have to plug in the point that it passes through, so that you can find b. For the perpendicular line: The slope is 1/9 and it passes through the point (6,3), so you have to plug in those values. \[3=(1/9)(6)+b\]\[\frac{27}{9}-\frac{6}{9}=b\]\[b=\frac{21}{9}\] Then for the parallel line: The slope is still -9 and it also passes through the point (6,3). \[3=(-9)(6)+b\]\[3=-27+b\]\[b=30\] Can you write the equations from this?

OpenStudy (anonymous):

y=1/9x + 21/9 y= -9x + 30?? is that right? @trainwrecking

OpenStudy (anonymous):

@trainwrecking ??

OpenStudy (anonymous):

Yeah, that's right! Good job!

OpenStudy (anonymous):

would y=-9x+54 work also? @trainwrecking

OpenStudy (anonymous):

or is it b=57?

OpenStudy (anonymous):

@rezende914 If you want to check to see if it works, just plug in the point and see if both sides end up equal. 3=-9(6)+54 That wouldn't work because it would evaluate as 3=-54+54. *Actually the answer I gave earlier was wrong. I'm sorry, I multiplied wrong! This is what should really happen: 3=(−9)(6)+b 3=-54+b b=57 Sorry! >.<